# ProFL: Performative Robust Optimal Federated Learning

### v1

@@ C.4 Additional Lemmas and their Proofs
The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemmas 5.8 and 5.9. $W_{1}\left(D,Q\right)_{\max}$ is the largest $W_{1}$ distance between $D_{i}(\theta)$ and $Q_{i}$ for all clients $i\in\mathcal{V}$ among all iteration $t\in[0,T-1]$.
Proof of Lemma C.2.
@@
The proof of the Lemma 12 of Izzo et al., (2021) shows that
@@
because the estimation error of $\nabla\mathcal{L}_{i,1}$ is much smaller than $\nabla\mathcal{L}_{i,2}$ by using Lemmas 5.8 and 5.9. Apply the fact that $\epsilon_{i}\in[0,1]$ and we can obtain
@@
Lemma C.3 (Sensitivity of contribution Dynamic).
For any client $i\in\mathcal{V}$, if Assumption 5.1and 5.3 hold, then with $D_{i}(\theta)=\nu_{i,1}(\theta)D_{i,1}(\theta)+\left(1-\nu_{i,1}(\theta)\right)D_{i,2}(\theta)$ for all $\theta,\theta^{\prime}\in\Theta$, we have:
@@
Appendix D Additional Experimental Results
D.1 Additional Experimental Results of Convergence to Performative Optimal Solution
Fig. 3 shows the results of case study house pricing regression, where an FL system is trained with 10 heterogeneous clients. Although PFL can stabilize the training process when data changes based on the model, it converges to an undesirable stable solution. In contrast, our algorithm converges to a stable solution with a much lower loss. In Fig. 3(a), the results show that both algorithms exhibit faster convergence rates with larger enrollment fraction. However, significant oscillations occur in PFL when the enrollment fraction is low, while such oscillations are much weaker in ProFL.
Fig. 3(b) shows the impact of varying degrees of heterogeneity, where $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$ with $\alpha\in\{0.1,0.3,1\}$. The results show that, despite the varying levels of heterogeneity, our algorithm maintains the capacity to converge to ${\theta}^{\text{PO}}$.
(a) Enrollment Fraction
(b) Different Heterogeneity
Figure 3: Linear Regression Results with Different Enrollment Fractions and Heterogeneity
(a) Sample Size
(b) Local Rounds $R$
Figure 4: Linear Regression Results with Different Sample Size and Local Rounds
In Fig. 4(a), the results show that both algorithms exhibit faster convergence rates with larger sample sizes. With a larger sample size, the results of PFL are closer to the performative stable point and exhibit fewer oscillations. However, PFL cannot converge to the performative optimal point, no matter how many samples are used. Fig. 4(b) demonstrates that a larger value of $R$ corresponds to faster convergence, reducing the communication time required in the FL system to reach convergence.
Fig. 5(a) shows the results of the case study pricing with dynamic contribution with 10 heterogeneous clients. The results indicate that, unlike PFL, which converges to $\theta^{\text{PS}}$, our algorithm converges to a solution, which is close to the performative optimal point. Fig. 5(b) shows the results of the case study regression with dynamic contribution with 10 heterogeneous clients. ProFL has faster convergence speed and less oscillation, although there is a gap between the results of both algorithms and the performative optimal point. The gap is a result of the non-linear $f_{i}(.)$, which has been discussed in more detail in the main paper.
(a) Pricing
(b) Regression
Figure 5: Contribution Dynamics
D.2 Performance when ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
In this section, we introduce a special case where $\theta^{\text{PO}}=\theta^{\text{PS}}$. This special case can arise in both strategic and dynamic contribution settings. We present the results in Fig. 6, demonstrating that when $\theta^{\text{PO}}=\theta^{\text{PS}}$, both algorithms converge to the same point.
(a) Strategic Dynamic
(b) Contribution Dynamic
Figure 6: Dynamics with ${\theta}^{\text{PO}}={\theta}^{\text{PS}}$
D.2.1 Performance without Model-dependent Shifts.
We also evaluate the algorithm when the data distribution does not change based on the model. In this case study, we use the Adult dataset for a binary classification task. Specifically, there are 10 homogeneous clients with $\gamma_{i,0}=\gamma_{i,1}=0$. We consider four features from the Adult dataset. When the data remains constant as the model is updated, both PFL and our algorithm reduce to FedAvg, and they converge to the same solution. This is illustrated in Fig. 6(a).
D.2.2 Performance when $\theta^{\text{PO}}=\theta^{\text{PS}}$ with Dynamic Contribution
Consider case study pricing with dynamic contribution with $K=2$ groups with different fixed distributions but a different one-dimensional fraction where $f_{i,1}(\theta)=0.5+0.5\theta$ and $f_{i,2}(\theta)=1-f_{i,1}(\theta)$. Under this special case, the performative stable and optimal points are identical, i.e., $\theta^{\text{PO}}=\theta^{\text{PS}}$. We can calculate $\theta^{\text{PO}}=\frac{a_{1}+a_{2}}{2(a_{1}-a_{2})}$ and $\theta^{\text{PS}}=\frac{a_{1}+a_{2}}{a_{1}-a_{2}}$. When $a_{1}+a_{2}=0$, $\theta^{\text{PO}}=\theta^{\text{PS}}$. In this case, we set $a_{1}=-a_{2}=0.5$. Fig. 6(b) illustrates the results for this specific case study, showing that both our algorithm and PFL converge to the same solution. However, our method exhibits faster convergence compared to PFL.
D.2.3 Figures of Table 2
The figures illustrating the experimental results from Table 2 in the main paper are shown in Fig. 7.
(a) Same Distributions
(b) Different Distributions
(c) Credit Dataset
(d) Adult Dataset
Figure 7: Binary Classification on Synthetic and Realistic Dataset
Appendix E Training details
@@
The experimental setup and parameter selections follow the methodology established in Izzo et al., (2021); Jin et al., (2024), ensuring consistency and comparability with existing results.
E.1 Realistic Data
@@
  • Adult data (Becker and Kohavi,, 1996) where the goal is to predict whether a person’s annual income exceeds \$50K based on 14 features (e.g., age, years of education).
@@
  • Give Me Some Credit data (Credit Fusion,, 2011) which has 11 features (e.g., monthly income, debt ratio) and can be utilized to predict whether a person has experienced a 90-day past-due delinquency or worse.
E.2 Loss Functions and Ridge Penalty
@@
E.3 Random Seeds
In our experiments, which show the results of 10 trials, the random seeds used range from 0 to 9.
E.4 Parameters and Distributions
The experimental parameters are listed in Table 3, and the distribution $D_{i}(\cdot)$ for each experiment is detailed in Table 4. The corresponding case studies for the experiments are shown in Table 5.
@@ E.4 Parameters and Distributions
Table 1: PG 0.01 -- 1 500 0
Table 1: ProFL 0.01 5 1 500 0
Table 2: Same $D_{i}$ 0.03 10 500 500 0
Table 2: Different $D_{i}$ 0.03 500 1 500 0
Table 2: Credit 0.01 5 1 5812 0
Table 2: Adult 0.03 5 1 2368 0
@@
Figure 3b PFL:0.005 ProFL:0.001 5 1 500 0
Figure 3a 0.03 5 1 2368 0
Figure 3b 0.03 5 1 2368 0
Figure 4a PFL:0.005 ProFL:0.001 5 1 5, 50, 500 0
Figure 4b PFL:0.005 ProFL:0.001 5,20,50 1 500 0
Figure 5a 0.03 5 1 500 0
Figure 5b 0.001 5 1 500 0
Figure 6a 0.03 5 1 500 0
Figure 6b 0.03 5 1 500 0
Table 4: Data Distributions in the Experiments
Figure/Table Distribution
Table 1: PG and ProFL $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0,0.25,0.5\}$
Table 2: Same $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$ and $X_{i}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$.
Table 2: Different $D_{i}$ $\gamma_{i,1}\in[2.8,3.2]$, $X_{1}\sim\mathcal{N}(-1-\gamma_{i,1}\theta,0.25)$
and $X_{2}\sim\mathcal{N}(-0.8-\gamma_{i,1}\theta,0.25)$ for $Y=1$;
$\gamma_{i,0}\in[0,0.04]$ and $X_{i}\sim\mathcal{N}(1-\gamma_{i,0}\theta,0.25)$ for $Y=0$
Table 2: Credit $\gamma_{i,1}=3$ and $\gamma_{i,0}=0.$ Give Me Some Credit datasetCredit Fusion, (2011)
Table 2: Adult $\gamma_{i,1}\in[2.8,3.2]$ and $\gamma_{i,0}\in[0,0.004].$ Adult dataset (Becker and Kohavi,, 1996)
Figure 1a $\mu_{0}\in[6,7]$, $\sigma=1$, $\gamma_{i}\in[1,3]$
Figure 1b,1c,2a-c $\gamma_{i}\in[0.09,0.11]$
Figure 3a $\gamma_{i}\in[1.5,1.8]$
Figure 3b $\gamma_{i}\in[1.5-\alpha,1.5+\alpha]$, $\alpha\in\{0.1,0.3,1\}$
Figure 4a $\gamma_{i}\in[1.5,1.8]$
Figure 4b $\gamma_{i}\in[1.5,1.8]$
Figure 5a $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$ $f(\theta)=0.5+0.5\theta$
Figure 5b $y=f_{i}(\theta)x+3(1-f_{i}(\theta))x+\varepsilon,\varepsilon\sim\mathcal{N}(0,4)$
Figure 6a $\gamma_{i}=0.$ 10 homogeneous clients with Adult dataset
Figure 6b $D_{i,1}\sim\mathcal{N}(0.5,0.25)$, $D_{i,2}\sim\mathcal{N}(-0.5,0.25)$
Table 5: Type of Case Study and Corresponding Figures or Tables
Figure/Table Type of Case Study
Table 1: house pricing regression
Table 2: binary strategic classification
Figure 1a pricing with dynamic demands
Figure 1b,1c,2a-c pricing with dynamic contribution
Figure 3a,3b,4a,4b house pricing regression
Figure 5a pricing with dynamic contribution
Figure 5b regression with dynamic contribution
Figure 6a binary strategic classification
Figure 6b pricing with dynamic contribution

### v2

@@ A.4 Additional Lemmas and their Proofs
The notation $(\boldsymbol{\omega_{D}})$ and $(\boldsymbol{\omega_{Q}})$ is defined in Lemmas 4.8 and 4.9. $W_{1}\left(D,Q\right)_{\max}$ is the largest $W_{1}$ distance between $D_{i}(\theta)$ and $Q_{i}$ for all clients $i\in\mathcal{V}$ among all iteration $t\in[0,T-1]$.
Proof of Lemma A.2.
@@
The proof of the Lemma 12 of pmlr-v139-izzo21a shows that
@@
because the estimation error of $\nabla\mathcal{L}_{i,1}$ is much smaller than $\nabla\mathcal{L}_{i,2}$ by using Lemmas 4.8 and 4.9. Apply the fact that $\epsilon_{i}\in[0,1]$ and we can obtain
@@
Lemma A.3 (Sensitivity of contribution Dynamic).
For any client $i\in\mathcal{V}$, if Assumption 4.1and 4.3 hold, then with $D_{i}(\theta)=\nu_{i,1}(\theta)D_{i,1}(\theta)+\left(1-\nu_{i,1}(\theta)\right)D_{i,2}(\theta)$ for all $\theta,\theta^{\prime}\in\Theta$, we have:
@@
Appendix B Case Studies
B.1 Pricing with Dynamic Demands
Consider a company that aims to find the proper prices for various goods. Let $\theta\in\mathbb{R}^{d}$ be the prices associated with the set of $d$ goods, and $Z_{i}\in\mathbb{R}^{d}$ be the retailer $i$’s demands for these goods. We assume $Z_{i}\sim\mathcal{N}(\mu(\theta),\sigma^{2})$ is model dependent Gaussian distribution, i.e., mean demand for each good decreases linearly as the price increases. $\mu(\theta)=\mu_{0}-\gamma_{i}\theta$, here $\gamma_{i}\geq 0$ measures the price sensitivity and varies across different retailers. The goal is to maximizes total revenue over all retailers $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{Z_{i}\sim D_{i}(\theta)}[\theta^{T}Z_{i}]$.
B.2 Pricing with Dynamic Contribution
Consider a similar setting as above where a company aims to find prices for various goods that maximize the total revenue. Suppose the retailers it faces have consumers from $K$ groups (e.g., old/young) and these consumers have fixed demands but they have options to purchase the goods from other companies. Here we assume the demands of retailer $i$’s consumers from group $k\in[K]$ have fixed distribution $D_{i,k}=\mathcal{N}(\mu_{i,k},\sigma_{i}^{2})$, but the fraction of groups change dynamically based on prices $\theta$. Let $r_{i,k}(\theta)=B-\mathbb{E}_{Z_{i}\sim D_{i,k}}[\theta^{T}Z_{i}]$ be the expected remaining budget of group $k$ after purchasing the goods (with initial budget $B$) and assume the fraction of group $k$ is $\nu_{i,k}(\theta)=\frac{r_{i,k}(\theta)}{\sum_{k=1}^{K}r_{i,k}(\theta)}$ (i.e., groups with more remaining budgets have more incentives to stay in the system). Then the retailer $i$’s total demands would follow the mixture of distributions $D_{i}(\theta)=\sum_{k=1}^{K}\nu_{i,k}(\theta)D_{i,k}$.
B.3 Binary Strategic Classification
[…]
B.4 House Pricing Regression
Consider a regression task that aims to find the proper listing prices for houses based on features such as size, age, number of bedrooms. Suppose the features of houses in district $i$ follow $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. Given predictive model $\theta$, let the listing price be $h(X_{i})=\theta^{T}X_{i}$. Since the houses with higher listing prices tend to lower the demand, the actual selling price $Y$ depends on $\theta$ and we assume $Y_{i}=(a_{i}-\gamma_{i}\theta)^{T}X_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}_{n})$ is Gaussian noise. The goal is to minimizes prediction error
B.5 Regression with dynamic contribution
Consider an example of retail inventory management, where the regression model $\theta$ predicts product sales considering features like seasonal trends and advertising. Distributor $i$ has retailers from two groups $k\in\{1,2\}$, and each retailer’s data distribution is fixed. The fraction of retailers from each group is $\nu_{i,k}(\theta)=\frac{\ell_{i,-k}(\theta)+c}{\sum_{k^{\prime}\in\{1,2\}}\left(\ell_{i,k^{\prime}}(\theta)+c\right)}$ where $k=\{1,2\}\setminus k$ because the retailers are more likely to stop purchasing if they experience higher prediction error $\ell_{i,k}(\theta)$. The goal is to find $\theta$ that minimizes expected prediction error $\sum_{i=1}^{N}\alpha_{i}\mathbb{E}_{(X_{i},Y_{i})\sim D_{i}(\theta)}\left[(\theta^{T}X_{i}-Y_{i})^{2}\right]$, where $Y_{i}$ is a weighted combination of fixed distributions.
Appendix C Training details
@@
The experimental setup and parameter selections follow the methodology established in pmlr-v139-izzo21a; jin2023performative, ensuring consistency and comparability with existing results.
C.1 Realistic Data
@@
  • Adult data misc_adult_2 where the goal is to predict whether a person’s annual income exceeds \$50K based on 14 features (e.g., age, years of education).
@@
  • Give Me Some Credit data GiveMeSomeCredit which has 11 features (e.g., monthly income, debt ratio) and can be utilized to predict whether a person has experienced a 90-day past-due delinquency or worse.
C.2 Loss Functions and Ridge Penalty
@@
C.3 Parameters and Distributions
The experimental parameters are listed in Table 2.
@@ C.3 Parameters and Distributions
Table I: PG 0.01 - 1 500 0
Table I: ProFL 0.01 5 1 500 0
@@
Figure 4a 0.03 10 500 500 0
Figure 4b 0.03 500 1 500 0
Figure 4c 0.01 5 1 5812 0
Figure 4d 0.03 5 1 2368 0
Figure 5a,5b PFL:0.005 ProFL:0.001 5 1 500 0
@@
Figure 6b PFL:0.005 ProFL:0.001 5 1 5, 50, 500 0
C.4 Experiments Compute Resources
CPU: 11th Gen Intel(R) Core(TM) i7-11800H @ 2.30GHz 2.30 GHz.